OpenStudy (anonymous):

if the length of each side of a triangle is cut to 1over3 of its original size, what happens to the area of the triangle? please help!!

OpenStudy (anonymous):

area will decrease by 3^2 or 1/9 th time of the original.

OpenStudy (anonymous):

if an equilateral triangle area= 1/2 l^2 sin 60 =\[\sqrt{3}/4\] l area original : 1/2 l^2 sin 60 area of the new one: 1/2 (l/3)^2 sin 60 1/2 (1/9)^2 sin 60 =\[\sqrt{3}/36 l\] \[\frac{ \sqrt{3}l/3}{ \sqrt{3}l/36 }\] 1/9th of the original